OpenStudy (anonymous):

y = 2x + 4 2x - y = 6 Solve the system of equations using substitution. A) (2,8) B) (2,-8) C) (10,24) D) no solution

OpenStudy (alexandervonhumboldt2):

y=2x+4 y=2x-6 2x-6=2x+4 0=- false no solution

OpenStudy (anonymous):

Thanks